OpenStudy (anonymous):

The cost for a package of five plain T-shirts is $11.35. If c represents the cost of each individual T-shirt, the equation 5c = 11.35 can be used to find the value of c. Solve for c.

OpenStudy (anonymous):

Since you're trying to find the cost of each T-shirt, the first thing to do is to write out your problem. 5c=11.35 Now we get c by itself and solve. 5c/5=11.35/5 c=2.27 Each T-shirt would cost $2.27.